The Engineering Manager in A Company Ltd. is responsible for maximising the profit from four product lines A, B, C and D. They have been provided with the following information Direct materials (kgs) 80 90 95 70 Direct labour (hrs) Welding Assembly Finishing 4.5 3.5 9.5 4 Selling price per unit (E) 1500 2500 3000 1000 Max.sales for the period in units) 400 300 280 450 250 250 250 250 Budgeted production (in units) Additional information is also provided i. There are no closing stocks ii. Material costs are estimated at £5.00/kg ii. The labour rate for the assembly and finishing operations is £15.00/hr and £20/hr for the welding operation iv. The skilled labour required for the welding process is in short supply although there is just enough to meet the current budgeted production (a) Prepare a statement showing the contribution from each product based on the budgeted production (8 marks) (b) Calculate the contribution per unit of limiting factor (2 marks) (c) Prepare a statement showing the maximum total contribution that can be obtained. Explain your rationale (10 marks)
Answer #1

Answer:

A)

Product A B C D
Direct Matrials  M 80 80 95 70
Direct labor
welding 4 4.5 4.5 3.5
Assembly 8 9 9.5 6
Finishing 4 4 4 3
Selling Price S 1500 2500 3000 1000
Max sales for period 400 300 280 450
Budgeted production P 250 250 250 250
Material cost MC=5*M 400 400 475 350
Labor cost
welding  WC=20*W 80 90 90 70
Assembly  AC=15*A 120 135 142.5 90
Finishing  FC=15*15 60 60 60 45
Contribution per unit C=S-MC-WC-AC-FC 840 1815 2232.5 445
Contribution for budgeted production B=P*C 210000 453750 558125 111250

B) Since welding labor are limited so Contribution per unit of welding labor are

Product A B C D
Direct labor
welding W 4 4.5 4.5 3.5
Budgeted production P 250 250 250 250
Total welding hours for budgeted production WH=P*W 1000 1125 1125 875
Contribution for budgeted production B =P*C 210000 453750 558125 111250
Contribution per welding hour = B/WH 210 403.33 496.11 127.14

C)

Since Skilled labor for welding process are enough for budgeted production so total skilled labor available are=1000+1125+1125+875=4125 labor hours

Since maximum contribution per unit of welding hour is for product C so we will produce maximum number product C. So total welding hours utilized =max sales for product C*4.5=280*4.5=1260 labor hours

Since second maximum contribution per unit of welding hour is for product B so we will produce maximum number product B. So total welding hours utilized =max sales for product B*4.5=300*4.5=1350 labor hours

Remaining labor hours=4125-1260-1350=1515 Labor hours

Since we are left with only 1515 labor units so we will produce third maximum contribution per unit of welding hour which is product A so production for product A=1515/4=378.75 units

So total contribution = Contribution from C+Contribution from B+Contribution from A

Contribution from C=280* contribution per unit of C=280*2232.5=625100

Contribution from B=300* contribution per unit of B=300*1815=544500

Contribution from A=378.75* contribution per unit of A=400*840=318150

total contribution =625100+544500+318150=1487750